OpenStudy (genny7):

Write an equation in slope-intercept form for the given situation. slope = 3/4, passes through (-6,9)

OpenStudy (anonymous):

y=mx+b m=3/4 x=-6 y=9 solve for b first

OpenStudy (anonymous):

use this